OpenStudy (justmeandme):

In ABCD ≅ MNOP, m∠A = x - 10 and m∠M =2x-30. What is m∠A? a. 10 b. 20 c. 30 d. 40

OpenStudy (fibonaccichick666):

so what does that weird looking equals sign mean?

OpenStudy (justmeandme):

congruent @FibonacciChick666

OpenStudy (fibonaccichick666):

good, so when two things are congruent, what does that mean?

OpenStudy (jiteshmeghwal9):

If ABCD is congruent to MNOP then angle A = angle m

OpenStudy (jiteshmeghwal9):

X-10=2x-30

OpenStudy (will.h):

Hint if <A is congruent to <m what does that mean for their measure of angle

OpenStudy (jiteshmeghwal9):

Find x

OpenStudy (jiteshmeghwal9):

Angle a =x-10

OpenStudy (will.h):

exactly find x and then once you get the value of x substitute that value in the x - 10 to find the measure of Angle A that's all

OpenStudy (will.h):

he/she obviosuly confused otherwise why would ask question so here's the full answer. we know that M<a = M<m that means the measure of Angle A is the same as the measure of angle m x - 10 = 2x - 30 solve for x 1st subtract 2x from both sides x - 2x - 10 = 2x - 2x - 30 -x - 10 = -30 now add 10 to both sides -x = -30 + 10 -x = -20 now divide -1 on both sides to get rid of the negative sign in x x = -20/-1 x = 20 Now that we found the value of x we may substitue it in the m<A = x - 10 to find the measure of angle A m<A = 20 - 10 m<A = 10 hope that helps NOTE: Since we said that m<A = m<M then that means there measures would be equal you can check if your answer correct or not by finding measure of angle m m<M = 2x - 30 m<M = 2(20) - 30 m<M = 40 - 30 m<M = 10 which is the same as m<A which means our answer is correct

OpenStudy (will.h):

let me know if you have any questions

OpenStudy (jiteshmeghwal9):

Perfect

OpenStudy (justmeandme):

Thank you so much @Will.H I was so confused

OpenStudy (will.h):

you welcome :)